# Lowering the Error Floor of Error Correction Code Transformer

### v1

@@ IV Proposed Loss Function for Hybrid Decoders
where $\mathbf{x}^{(i)}$ can be fixed as the zero codeword and $\mathbf{y}^{(i)}$ is sampled as $\mathbf{x}_{s}^{(i)}+\mathbf{z}^{(i)}$, retaining only the samples that satisfy $d_{H}(\mathbf{x}^{(i)},\mathbf{y}^{(i)})>t_{c}$. The noise variance $\sigma^{2}$ is uniformly sampled from the training SNR range. In addition, the step function $u(x)$ is implemented via a straight-through estimator (STE) [16] with the sigmoid function as a proxy, as described in [17]. The Hamming distance $d_{H}(\mathbf{x},\widetilde{f}_{\boldsymbol{\theta}}(\mathbf{y}))$ is estimated by multiplying the code length $n$ with the soft BER loss defined in [18].
@@ V Experimental Results
In this section, we present the experimental results for the proposed hybrid decoding algorithms. To evaluate its effectiveness, we measure the BER and FER performance for four BCH codes: $(31,16)$, $(63,36)$, $(63,45)$, and $(63,51)$, which have error correcting capabilities of $t=3$, $t=5$, $t=3$, and $t=2$, respectively. In all simulations involving ECCT, the number of decoding layers $N$ is fixed to $6$, and the embedding dimension $d$ is set to $128$. Additionally, the parity check matrices used in the experiments and maximum likelihood (ML) decoding results are obtained from [14].
Fig. 4 shows the FER performance of the proposed hybrid decoder along with other decoders: ML, hard decision, and the original ECCT. The hybrid decoder incorporates hard decision decoders as both pre- and post-decoders and is trained using the proposed loss (IV). The results indicate that the hybrid decoder effectively mitigates the error floor, significantly outperforming the original ECCT in the high SNR region. Furthermore, even in the waterfall region, the hybrid decoder outperforms both the original ECCT and hard decision decoder. Notably, the hybrid decoder achieves the ML decoding performance for the $(31,16)$ BCH code. By combining two distinct decoders and training them with the proposed loss, the hybrid decoding approach can achieve significant coding gains.
TABLE I: Comparison of FER of Decoders with Different Configurations
Methods (31, 16) BCH code (63, 45) BCH code
[…]
ECCT [4] $1.52\mathrm{e}{-2}$ $2.53\mathrm{e}{-3}$ $2.86\mathrm{e}{-4}$ $6.41\mathrm{e}{-2}$ $8.65\mathrm{e}{-3}$ $4.68\mathrm{e}{-4}$
Pre + ECCT $1.45\mathrm{e}{-2}$ $2.43\mathrm{e}{-3}$ $2.76\mathrm{e}{-4}$ $4.96\mathrm{e}{-2}$ $5.63\mathrm{e}{-3}$ $2.77\mathrm{e}{-4}$
ECCT + Post $6.70\mathrm{e}{-3}$ $9.23\mathrm{e}{-4}$ $7.47\mathrm{e}{-5}$ $3.36\mathrm{e}{-2}$ $3.42\mathrm{e}{-3}$ $1.26\mathrm{e}{-4}$
Pre + ECCT + Post $6.47\mathrm{e}{-3}$ $8.96\mathrm{e}{-4}$ $7.25\mathrm{e}{-5}$ $2.91\mathrm{e}{-2}$ $2.55\mathrm{e}{-3}$ $8.63\mathrm{e}{-5}$
Pre + ECCT + Post + loss (IV) $\mathbf{2.23\mathrm{\mathbf{e}}{-3}}$ $\mathbf{1.63\mathrm{\mathbf{e}}{-4}}$ $\mathbf{5.41\mathrm{\mathbf{e}}{-6}}$ $\mathbf{2.26\mathrm{\mathbf{e}}{-2}}$ $\mathbf{1.56\mathrm{\mathbf{e}}{-3}}$ $\mathbf{4.26\mathrm{\mathbf{e}}{-5}}$
To assess the impact of each contribution within the hybrid decoding framework, we compare the FER performance of different decoder configurations, as shown in Table I. In this table, “Pre” and “Post” represent hard decision pre- and post-decoders, respectively while “ECCT” refers to the original ECCT trained on the binary cross entropy loss function (II-B). The results show that the hybrid decoder combining ECCT with both pre- and post-decoders outperforms decoder configurations using only a single auxiliary decoder, implying the importance of both pre- and post-decoders. Moreover, the hybrid decoder trained with the proposed loss (IV) achieves the best FER performance among all configurations. This confirms that the proposed loss function effectively improves ECCT within the hybrid decoding framework, thereby enhancing overall decoding performance.
@@ VI Conclusion
In this paper, we presented the first observation of the error floor problem of ECCT. To address this issue, we proposed hybrid decoding algorithms for ECCT, along with an optimized loss function. By integrating ECCT with hard decision decoders, the hybrid decoding approach effectively corrects a wider range of errors, thereby enhancing decoding performance. The proposed loss function enables ECCT to focus on correcting only critical errors, optimizing the overall hybrid decoding process. Our simulation results show that the hybrid decoder effectively lowers the error floor, achieving improved FER performance in the high SNR region. Moreover, the hybrid decoder significantly enhances decoding performance in the waterfall region. The proposed hybrid decoding algorithm and its associated training method, which accounts for the pre- and post-decoders, can be generalized to other neural decoders. This approach provides a versatile strategy for integrating neural decoders and conventional decoders to effectively lower the error floor and enhance decoding performance.
@@ Acknowledgment
This work was partly supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (RS-2024-00398449, Network Research Center: Advanced Channel Coding and Channel Estimation Technologies for Wireless Communication Evolution) and the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2023-00212103).

### v2

@@ IV Proposed Loss Function for Hybrid Decoders
where $\mathbf{x}^{(i)}$ can be fixed as the zero codeword and $\mathbf{y}^{(i)}$ is sampled as $\mathbf{x}_{s}^{(i)}+\mathbf{z}^{(i)}$, retaining only the samples that satisfy $d_{H}(\mathbf{x}^{(i)},\tilde{\mathbf{y}}^{(i)})>t_{c}$. Note that, to simplify the sampling process, we simply check whether $d_{H}(\mathbf{x}^{(i)},\tilde{\mathbf{y}}^{(i)})>t_{c}$, instead of checking the syndrome of the hard decision decoder by executing a hard decision decoder such as the Berlekamp-Massey algorithm [23] for every sample. The noise variance $\sigma^{2}$ is uniformly sampled from the training SNR range. In addition, the step function $u(\alpha)$ is implemented via a straight-through estimator (STE) [24] with the sigmoid function as a proxy, as described in [25]. The Hamming distance $d_{H}(\mathbf{x},\widetilde{f}_{\boldsymbol{\theta}}(\mathbf{y}))$ is estimated by multiplying the code length $n$ with the soft BER loss defined in [26]. In the case of the hybrid decoder architecture with only a pre-decoder, as shown in Fig. 2LABEL:sub@fig:architecture_pre, the loss is computed by removing the step function from (IV). For the hybrid decoder with only a post-decoder, as in Fig. 2LABEL:sub@fig:architecture_post, the condition $d_{H}(\mathbf{x}^{(i)},\mathbf{y}^{(i)})>t_{c}$ is no longer checked.
TABLE I: Comparison of FER of Decoders with Different Configurations in Error Floor Region
Methods (31, 16) BCH code (63, 36) BCH code
[…]
ECCT [4] $2.54\mathrm{e}{-4}$ $1.94\mathrm{e}{-5}$ $9.42\mathrm{e}{-7}$ $1.59\mathrm{e}{-3}$ $7.44\mathrm{e}{-5}$ $1.35\mathrm{e}{-6}$
Pre + ECCT $8.45\mathrm{e}{-4}$ $9.23\mathrm{e}{-5}$ $4.76\mathrm{e}{-6}$ $2.53\mathrm{e}{-4}$ $5.08\mathrm{e}{-6}$ $3.26\mathrm{e}{-8}$
ECCT + Post $7.78\mathrm{e}{-5}$ $4.25\mathrm{e}{-6}$ $1.45\mathrm{e}{-7}$ $1.42\mathrm{e}{-4}$ $2.51\mathrm{e}{-6}$ $1.76\mathrm{e}{-8}$
Pre + ECCT + Post $6.95\mathrm{e}{-4}$ $7.59\mathrm{e}{-5}$ $4.20\mathrm{e}{-6}$ $7.69\mathrm{e}{-5}$ $9.48\mathrm{e}{-7}$ $8.29\mathrm{e}{-9}$
ECCT + Post + loss (IV) $\mathbf{7.32\mathrm{\mathbf{e}}{-5}}$ $\mathbf{2.38\mathrm{\mathbf{e}}{-6}}$ $\mathbf{5.02\mathrm{\mathbf{e}}{-8}}$
Pre + ECCT + Post + loss (IV) $\mathbf{7.00\mathrm{\mathbf{e}}{-5}}$ $\mathbf{6.60\mathrm{\mathbf{e}}{-7}}$ $\mathbf{6.38\mathrm{\mathbf{e}}{-9}}$
@@ V Experimental Results
In this section, we present the experimental results for the hybrid decoding scheme, where the ECCT decoder is trained with the proposed loss function. To evaluate its effectiveness, we measure the FER performance for four BCH codes: $(31,16)$, $(63,36)$, $(63,45)$, and $(63,51)$, which have error correcting capabilities of $t_{c}=3$, $t_{c}=5$, $t_{c}=3$, and $t_{c}=2$, respectively. In all simulations involving ECCT, the number of decoding layers $N$ is fixed to $6$, and the embedding dimension $d$ is set to $128$. Furthermore, the parity check matrices used in the experiments and maximum likelihood (ML) decoding results are obtained from [20].
Fig. 4 shows the FER performance of the proposed hybrid decoder along with other decoders: ML decoding, hard decision decoding, and the original ECCT. The hybrid decoder combining both pre- and post-decoders is used for $(63,36)$ BCH code, while the hybrid decoder with only the post-decoder is used for $(31,16)$, $(63,45)$, and $(63,51)$ BCH codes, which yielded the best results. ECCTs in each hybrid decoder are trained using the proposed loss (IV), with necessary adjustments based on the hybrid decoder architecture. The results indicate that the hybrid decoder effectively mitigates the error floor, significantly outperforming the original ECCT in the high SNR region. Furthermore, even in the waterfall region, the hybrid decoder outperforms both the original ECCT and the hard decision decoder. By combining two distinct decoders and training them with the proposed loss, the hybrid decoding approach can achieve significant coding gains.
To assess the impact of each component in the hybrid decoding framework, we compare the FER performance of different decoder configurations, as shown in Table I. In this table, “Pre” and “Post” represent hard decision pre- and post-decoders, respectively while “ECCT” refers to the original ECCT trained on the binary cross entropy loss (II-B). Among the various decoder architectures, the hybrid decoder with only a post-decoder yields the best performance for the $(31,16)$ BCH code, whereas the architecture with both pre- and post-decoders performs best for the $(63,36)$ BCH code. This result is attributed to the higher undetected error rate for short codes. Furthermore, for both codes, training the hybrid decoder with the proposed loss (IV) further improves FER performance compared to training with the binary cross entropy loss (II-B).
@@ VI Conclusion
In this paper, we presented the first observation of the error floor phenomenon in ECCT. To address this issue, we adopted hybrid decoding strategy that integrates ECCT with hard decision decoders, combined with the proposed loss function tailored to this architecture. The hybrid decoding approach effectively broadens the range of correctable errors, thereby mitigating the error floor. In particular, the proposed loss function directs ECCT to focus on correcting critical errors, improving the effectiveness of the hybrid decoding process. Our simulation results support that the hybrid decoders trained with the proposed loss function can effectively lower the error floor, achieving improved FER performance in the high SNR region. Additionally, the hybrid decoder significantly improves decoding performance even in the waterfall region. The hybrid decoding strategy, along with its training method that accounts for both pre- and post-decoders, can be generalized to other neural decoders. This framework provides a flexible strategy for integrating neural decoders and conventional decoders to lower the error floor and improve overall decoding performance. In future work, exploring hybrid decoding strategies that combine ECCT with conventional BP decoders would be both valuable and promising.
@@ Acknowledgment
This work was partly supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korean government (MSIT) (RS-2024-00398449, Network Research Center: Advanced Channel Coding and Channel Estimation Technologies for Wireless Communication Evolution) and the National Research Foundation of Korea (NRF) grant funded by the Korean government (MSIT) (No. RS-2023-00212103).